to the previous experience, only 20% of the elderly will welcome the visit while 80% of the elderly will reject the visit. Mary, one of the part-time interviewers, will be responsible to visit 13 elderly residents today. Use X to denote the number of successful visits she will make today.
(b) Most likely, how many elderly residents will welcome Mary’s visit? Calculate the corresponding probability.
(c) Suppose Mary will spend 5 minutes in each rejected visit and 30 minutes in each successful visit. Use Y to denote the total time Mary will spend today. Express Y in terms of X. Hence, find the (i)minimum, (ii) maximum, and (iii) expected total time (in minutes) for Mary to finish all visits today?
(d) Mary will get $30 allowance for each rejected visit and $100 allowance for each successful visit. UseW to denote the total allowance Mary will get today. Express W in terms of X. Hence, find the (i)minimum, (ii) maximum, and (iii) expected total allowance Mary will get today?
1
Expert's answer
2019-11-12T12:13:19-0500

(b) "E(X)=pn=0.2*13=2.6"

The most likely outcome satisfies following inequality.

"np-q \leq k \leq np+p"

"13 \cdot 0.2-0.8 \leq k \leq 13 \cdot 0.2+0.2"

"1.8 \leq k \leq 2.8"

Thus, most likely, 2 elderly residents will welcome Mary’s visit.

The corresponding probability: "P(X=2)=C_{13}^2(0.2)^2(1-0.2)^{11}=0.2680"


(c) "Y=30X+5(13-X)=25X+65"

(i) "Y_{min}=25*0+65=65\;min"

(ii) "Y_{max}=25*13+65=390\;min"

(iii) "Y_{exp}=25*2.6+65=130\;min"


(d) "W=100x+30(13-x)=70X+390"

(i) "W_{min}=70*0+390=390"

(ii) "W_{max}=70*13+390=1300"

(iii) "W_{exp}=70*2.6+390=572"